Solve for x: (this will be an inverse trigonometric expression, not a numeric answer). 25y=5cos(3x-6)
Answer by Theo(13342)

i ran out of time but i think this is the correct answer.
check it out and see if you agree.
i graphed it and it looks ok but i don't have the time to extensively test it, although i'm pretty sure it's what you need.

basic definition is that y = cos(x) if and only if x = arc-cosine(y)

in your problem, you would use this as follows:

start with 25y = 5 * cosine(3x-6)

divide both sides of the equation by 5 to get:

5y = cosine(3x-6)

this is true if and only if arc-cosine(5y) = 3x - 6

add 6 to both sides of the equation to get:

arc-cosine(5y) + 6 = 3x

divide both sides of the equation by 3 to get:

x = (arc-cosine(5y) + 6) / 3

that should be your solution.

i tested it on 5y = 1/2

if 5y = 1/2, then arc-cosine (5y) = 60 degrees.

can't continue.

i have to leave now.

if you have further questions on this send them to me and i'll look at them later this evening or tomorrow morning.